Bad Debts, Accounts Receivable, etc...

At December 31, 2011, Ethan Company reports the following results for its calendar year.

Cash sales.... $1,803,750
Credit sales....$3,534,000

In addition, its unadjusted trial balance includes the following items.

Accounts receivable..... $1,070,100 debit
Allowance for doubtful accounts.... $15,750 debit

1. Prepare the following entry for this company to recognize bad debts under each of the following independent assumptions.
a) Bad debts are estimated to be 2% of credit sales.
b) An aging analysis estimates that 5% of year-end accounts receivable are uncollectible.

2. Show how Accounts Receivable and the Allowance for Doubtful Accounts appear on its December 31, 2011, balance sheet given the facts in part 1a.

3. Show how Accounts Receivable and the Allowance for Doubtful Accounts appear on its December 31, 2011, balance sheet given the facts in part 1b.
Answer #1

1. Prepare the following entry for this company to recognize bad debts under each of the following independent assumptions.

a) Bad debts are estimated to be 2% of credit sales.

2% of credit sales = 0.02*3,534,000 = 70,680

Debit Bad Debt Expense 70,680

Credit Allowance for doubtful accounts 70,680

b) An aging analysis estimates that 5% of year-end accounts receivable are uncollectible.

5% of year-end accounts receivable = 0.05*1,070,100 = 53,505

There is currently debit balance in the allowance account of 15,750, and the company needs it to be a credit balance of 53,505. So they need toincrease the allowance account by 15,750 + 53,505 = 69,255

Debit Bad Debt Expense 69,255

Credit Allowance for Doubtful Accounts 69,255

2. Show how Accounts Receivable and the Allowance for Doubtful Accounts appear on its December 31, 2011, balance sheet given the facts in part 1a.

Current Assets:

Accounts receivable…………………..…1,070,100

Less Allowance for doubtful accounts..54,930…….1,015,170

3. Show how Accounts Receivable and the Allowance for Doubtful Accounts appear on its December 31, 2011, balance sheet given the facts in part 1b.

Current Assets:

Accounts receivable…………………..…1,070,100

Less Allowance for doubtful accounts..53,505…….1,016,595

answered by: Ellie